# PropNet: a White-Box and Human-Like Network for Sentence Representation

### v1

@@ 4.1 Comparison with Embedding Models
Implementation Details. The experiment employs a comparison module which consists of two components, (1) the calculation of the difference vector and (2) prediction using CART models. Figure 3 presents the details of this module. After computing the difference vectors using the approach introduced in Section 3.5 for all pairs within the benchmark datesets, two CART models are trained on these vectors: one for short pairs P1-/P2 and another for long pairs P3+. These models are denoted as CART(P1-/P2) and CART(P3+) repectively. We call the API provided by scikit-learn  [^4] API: https://scikit-learn.org/stable/modules/generated/sklearn.tree.DecisionTreeRegressor.html for training. Since STS 2012-2016 and STS-B share the same rules for ground scores, training is implemented only on the train set of STS-B to learn these rules. Similarly, for SICK-R, two CART models are trained on its training set.
Predictions for short pairs and long pairs in the test set are concatenated before computing Spearman’s correlation. The random seed is fixed as 0, and the minimum number of samples at a leaf node is set to 10 for all trainings.
Results. In Table 5, a significant gap is observed between PropNet and the latest embedding model, AoE, which scores above 80. This indicates the efficacy of big-data-trained embeddings in comprehending text semantics. This table also demonstrates that embedding models have undergone notable development over the last ten years, with scores ranging from 65.01 to 82.43. Since PropNet marks the inception of a completely new research direction, offering purely white-box and human-like sentence representation, we are confident that PropNet will improve in the future, considering that there is substantial room for enhancement in its current version.
@@ 4.2 Investigation of Genres
Dataset. We utilize the STS-B test set  [^5] Approximately 8.7% of the STS-B records lack similarity scores due to dependency parsing errors. These records are excluded from the analysis. along with its predicted scores for analysis. The test set is split into three parts according to its field genre, which has three values: main-captions, main-news and main-forums. They represent image/video description, news and forum question/answer, respectively. Appendix C.1 provides more detailed explanations of these genres.
@@
For main-forums, PropNet currently lacks nodes to measure the likelihood of an action, such as modal verbs (e.g., "can", "should") and negation indicators (e.g., "not", "hardly"). In addition, main-forums also suffers from the coarseness of the evolutionary graph. For instance, expressions like "It is up to you" cannot be represented properly.
@@ 4.3 Human Cognitive Differences in STS Tasks
This section studies the cognitive process underlying how humans assign a score according to task instructions when given a sentence pair. Intuitively, within human cognition, each verb-related dimension adopted in Section 3.3 ought to contribute differently to the disparity between two sentences. In addition, for the same dimension, human perceptions of the disparity should vary to a certain extent. We aim to quantify these cognitive differences in this experiment.
Dataset. The P1- pairs of the STS-B dataset, whose sentences differ at only one dimension, are selected. Compared to the black-box nature of embedding models, PropNet significantly facilitates this selection process by simply checking if the corresponding element of difference vector is 2. Only pairs from main-captions and main-news are used, since main-forums contains too few samples. In total, 133 valid pairs are collected for analysis. All pairs and their distribution are detailed in Appendix C.2.1.
@@ 5 Discussion
Long Text Understanding. Although in this work we mainly discuss the representation of a single sentence, it is possible to extend this method to a discourse or a conversation. More sophisticated merging strategies are required to link the representations of each sentence into a unified network. This will provide novel perspectives for the research of long-text understanding in a white-box and human-like manner.
Cognitive Process Analysis. With the emergence of deep learning algorithms, the past decade has witnessed the generation of massive amounts of manually labeled data for training models. Therefore, cognitive processes underlying how humans label input data according to specific instructions, are recorded in these vast datasets. White-box and human-like approaches have the potential to illuminate the cognitive processes, which is extremely beneficial to gain a deeper understanding of humanity, such as our behaviors, mentality, and intelligence.
@@ 6 Limitation
PropNet lacks the ability to handle logic relations among propositions of a sentence. Evolutionary nodes of logic module would be added in future work. In addition, PropNet would fail to represent if there are errors from spaCy part-of-speech tagging and dependency parsing results. For instance, in the sentence “Blizzard wallops US Northeast, closes road”, the subject of “close” is identified as “wallop”.
@@ B.2 Similarity between Two Words
Given two words, we use Doubao  [^6] https://team.doubao.com/en/direction/llm to compute a similarity score ranging from 0 to 1. Since this task is straightforward, other LLM models like ChatGPT or Llama should work as well.
@@ B.3 Dimensional Similarity Code
Given two words, a code of 0 means two words are completely or almost the same. A code of 1 means they are similar. Note that when we compare "unknown" with one single word, they are regarded as similar. For example, in the comparison of "I cut an onion with a knife" and "I cut an onion", the dimension #action|#aux_obj has "knife" for the first sentence and "unknown" for the second sentence, and this comparison receives a code of 1. A code of 2 means the two words are completely different. LLM similarity scores along with thresholds are used to assign these DSC codes.
@@
Finally, given two groups of words with an equal length 2, we consider a binary group similarity score, which is then used to assign DSC codes 0, 1, and 2 using 0.2 and 0.7 as the thresholds. The binary group similarity score is calculated by the following steps:
@@
    If each group contains only one remaining word, calculate the similarity score between these two words and add it to score.
@@ Appendix C Experiments
C.1 Genres
@@
C.2 Cognitive Differences in STS Tasks
C.2.1 Experimental Samples
We select the P1- pairs from the STS-B dataset, which consist of sentences differing in only one dimension. Eight dimensions related to the verb of a proposition are considered: #action, #subject, #object, #where, #aux_obj, #goal, #source, and #reason. For example, "A man is speaking" and "A man is cooking" is a valid pair for the #action dimension, since they differ only in #action.
@@ C.2.2 Statistical Testing
C.2.3 Ground Score Rules

### v3

@@ 4.1 Comparison with Embedding Models
Implementation Details. The experiment employs a comparison module which consists of two components, (1) the calculation of the difference vector and (2) prediction using Classification and Regression Trees (CART) [12]. Figure 4 presents the details of this module. After computing the difference vectors using the approach introduced in Section 3.5 for all pairs within the benchmark datasets, two CART models are trained on these vectors: one for short pairs P1-/P2 and another for long pairs P3+, because the sizes of the difference vectors are different. These models are denoted as CART(P1-/P2) and CART(P3+) respectively. We call the API provided by scikit-learn  [^4] API: https://scikit-learn.org/stable/modules/generated/sklearn.tree.DecisionTreeRegressor.html for training. Since STS 2012-2016 and STS-B share the same rules for ground scores, training is implemented only on the train set of STS-B to learn these rules. Similarly, for SICK-R, two CART models are trained on its training set. The random seed is fixed as 0, and the minimum number of samples at a leaf node is set to 10. Further training details are appended in Appendix C.1.
Note that the predictions for short pairs and long pairs in the test set are concatenated before computing the Spearman’s correlation.
Results. In Table 5, a significant gap is observed between PropNet and the latest embedding model, AoE, which scores above 80. This indicates the efficacy of big-data-trained embeddings in comprehending text semantics. This table also demonstrates that embedding models have undergone notable development over the last ten years, with scores ranging from 65.01 to 82.43. Since PropNet marks the inception of a completely new research direction, offering purely white-box and human-like sentence representation, we are confident that PropNet will improve in the future, considering that there is substantial room for enhancement in its current version, which is presented in the next section.
@@ 4.2 Investigation of Genres
Dataset. We utilize the STS-B test set  [^5] Approximately 8.7% of the STS-B records lack similarity scores due to spaCy dependency parsing errors. These records are excluded from the analysis. along with its predicted scores for analysis. The test set is split into three parts according to its field genre, which has three values: main-captions, main-news and main-forums. They represent image/video description, news and forum question/answer, respectively. Appendix C.2 provides more detailed explanations of these genres.
@@
For main-forums, PropNet currently lacks nodes to measure the likelihood of an action, such as modal verbs (e.g., "can", "should") and negation indicators (e.g., "not", "hardly"). For example, it fails in comparing “You should do it” and “You can do it”, or “It’s not a good idea” and “It’s a good question”. In addition, main-forums also suffers from the coarseness of the evolutionary graph. For instance, expressions like "It is up to you" cannot be represented properly.
@@ 4.3 Human Cognitive Differences in STS Tasks
This section studies the cognitive process underlying how humans assign a score according to task instructions when given a sentence pair. Intuitively, different dimensions ought to contribute differently to the disparity between two sentences. In addition, for the same dimension, human perceptions of the disparity should vary to a certain extent. We aim to quantify these cognitive differences in this experiment.
Eight dimensions from Section 3.5.1 are considered: #action, #subject, #object, #where, #aux_obj, #goal, #source, and #reason. We aggregate #aux_obj, #goal, #source, and #reason into one dimension, labeled as #other.
Dataset. The P1- pairs of the STS-B dataset, whose sentences differ at only one dimension, are selected. Compared to the black-box nature of embedding models, PropNet significantly facilitates this selection process by simply checking if the corresponding element of difference vector is 2. For example, "A man is speaking" and "A man is cooking" is a valid pair for the #action dimension, since the difference vector has a value 2 at #action and 0 at all other positions. Only pairs from main-captions and main-news are used, since main-forums contains too few samples. In total, 133 valid pairs are collected for analysis. All pairs and their distribution are detailed in Appendix C.3.1.
@@ 5 Discussion
Long Text Understanding. Although in this work we mainly discuss the representation of a single sentence, it is possible to extend this method to a discourse or a conversation. Instance nodes enable PropNet to have potentials in addressing coreference resolution and representing contextual information effectively. Minor changes to current merging strategies can meet the need for linking the representations of each sentence into a unified network. This will provide novel perspectives for the research of long-text understanding in a white-box and human-like manner.
Cognitive Process Analysis. With the emergence of deep learning algorithms, the past decade has witnessed the generation of massive amounts of manually labeled data for training models. Therefore, cognitive processes underlying how humans label input data according to specific instructions, are recorded in these vast datasets. White-box and human-like approaches have the potential to illuminate the cognitive processes, which is exemplified by Section 4.3. It can significantly contribute to a deeper understanding of humanity, such as our behaviors, mentality, and intelligence.
@@ 6 Limitation
PropNet would fail to represent if there are errors from spaCy part-of-speech tagging and dependency parsing results. For instance, in the sentence “Blizzard wallops US Northeast, closes road”, the subject of “close” is identified as “wallop”.
@@ B.2 Similarity between Two Words
Given two words, we use Doubao  [^6] https://team.doubao.com/en/direction/llm to compute a similarity score ranging from 0 to 1. Doubao has an extremely low inference cost, while offering top-tier inference performance and an easy-to-use API for calling. The quality of the output scores are validated by manual inspection during development. Since this task is straightforward, other LLM models like ChatGPT or Llama should work as well.
@@ B.3 Dimensional Similarity Code
Given two words, a code of 0 means two words are completely or almost the same. A code of 1 means they are similar, sharing some important details. Note that when we compare "unknown" with one single word or a group of words, they are regarded as similar. For example, in the comparison of "I cut an onion with a knife" and "I cut an onion", the dimension #action|#aux_obj has "knife" for the first sentence and "unknown" for the second sentence, and this comparison receives a code of 1. A code of 2 means the two words are completely different. We manually check word pair cases along with their LLM scores to determine proper thresholds, best matching the DSC code standard. If LLM similarity score is no less than 0.7, a code of 0 is assigned. If LLM similarity score is less than 0.2, a code of 2 is assigned. Otherwise, a code of 1 is assigned.
@@
Finally, given two groups of words with an equal length 2, we consider a binary group similarity score which is calculated by the following steps:
@@
    If each group contains only one remaining word, calculate the similarity score between these two words and add it to score. Otherwise do nothing.
@@
This process is exemplified step by step by comparing ["apple", "banana"] and ["apple", "orange"] as follows:
  1. (1)
    Initialize score as 0.
  2. (2)
    Intersection is ["apple"] and its length is 1.
  3. (3)
    score = 1 + score. Now score is 1.
  4. (4)
    Remove "apple" from both groups. Now they become ["banana"] and ["orange"].
  5. (5)
    Since each group has only one word left, LLM is used to calculate the similarity score which outputs a 0.2. Now score is 1.2.
  6. (6)
    Binary group similarity score = score / 2 = 0.6
As the final score lies between 0.2 and 0.7, a DSC code 1 (similar) is assigned for this binary-word groups comparison.
@@ Appendix C Experiments
C.1 CART Training Details for STS Tasks
For the STS-B dataset, we tune only one parameter for CART training, the minimum number of samples at a leaf node (short as min_samples_leaf), on its train.jsonl (5749 samples) and validation.jsonl (1500 samples). The default values are adopted for other CART parameters, since finetuning models for higher Spearman’s score is not a key point in our work. What we place high weight on, is finding the weaknesses of PropNet and its opitimization directions in future research.
The SICK-R dataset contains a field named SemEval_set, indicating the train/dev/test split. We use the training split (4439 samples) and dev split (495 samples) for tuning min_samples_leaf.
CART Training is supported by the Scikit-Learn APIs, whose version is 1.6.0.
C.2 Genres
@@
C.3 Cognitive Differences in STS Tasks
C.3.1 Experimental Samples
@@ C.3.2 Statistical Testing
C.3.3 Ground Score Rules